# neurips_data_2024.bib

### v1

@@ Abstract
Dynamic link prediction is an important problem considered by many recent works proposing various approaches for learning temporal edge patterns. To assess their efficacy, models are evaluated on publicly available benchmark datasets involving continuous-time and discrete-time temporal graphs. However, as we show in this work, the suitability of common batch-oriented evaluation depends on the datasets’ characteristics, which can cause two issues: First, for continuous-time temporal graphs, fixed-size batches create time windows with different durations, resulting in an inconsistent dynamic link prediction task. Second, for discrete-time temporal graphs, the sequence of batches can additionally introduce temporal dependencies that are not present in the data. In this work, we empirically show that this common evaluation approach leads to skewed model performance and hinders the fair comparison of methods. We mitigate this problem by reformulating dynamic link prediction as a link forecasting task that better accounts for temporal information present in the data. We provide implementations of our new evaluation method for commonly used graph learning frameworks.
@@ 1 Introduction
Many scientific fields study data that can be modeled as graphs, where nodes represent entities that are connected by edges. Examples include social [ComputationalSocialScience2009lazer], financial [PhysicsFinancialNetworks2021bardoscia], biological [GenomicRegulatoryNetwork2002davidson] as well as molecular networks [david2020molecular]. Apart from the mere topology of interactions, i.e., who is connected to whom, such data increasingly include information on when these interactions occur. Depending on the temporal resolution, the resulting temporal graphs are often categorized as continuous-time or discrete-time [GraphNeuralNetworks2023longa]: State-of-the-art data collection technology provides high-resolution continuous-time temporal graphs, which capture the exact (and possibly unique) occurrence time of each interaction. Examples include time-stamped online interactions [PredictingDynamicEmbedding2019kumar] or social networks captured via high-resolution proximity sensing technologies [EstimatingPotentialInfection2013vanhems]. In contrast, discrete-time temporal graphs give rise to a temporally ordered sequence of static snapshots, where each snapshot contains interactions recorded within a (typically coarse-grained) time interval. Examples include scholarly collaboration or citation graphs, which frequently include monthly or yearly snapshots.
Building on the growing importance of temporal data and the success of graph neural networks (GNNs) for static graphs [GeometricDeepLearning2017bronstein, GraphNeuralNetworks2024corso], deep graph learning has recently been extended to temporal (or dynamic) graphs [feng2024comprehensive]. To this end, several temporal graph neural network (TGNN) architectures have been proposed that are able to simultaneously learn temporal and topological patterns. These architectures are often evaluated in dynamic link prediction, where the task is to predict the existence of edges in a future time window of length $\Delta t$, e.g., to provide recommendations to users [PredictingDynamicEmbedding2019kumar].
For dynamic link prediction, TGNNs commonly utilize temporal batches to speed up training [su2024pres]. To construct these temporal batches, the sequence of temporally ordered edges is divided into a sequence of equally large chunks that contain the same number of edges. Within each batch, edges are typically treated as if they occurred simultaneously, thus discarding temporal information within a batch. For continuous-time temporal graphs, such fixed-size batches are also likely to be associated with time windows of varying lengths $\Delta t$. Changing the batch size affects the resulting window lengths and could, e.g., change the task from predicting at the minute to the hour level, thus altering its difficulty (Figure 1a). In discrete-time temporal graphs, snapshots are typically so large that they comprise multiple batches (Figure 1b). This leads to the issue that the ordered sequence of batches does not necessarily correspond to a temporally ordered sequence. A batch-wise training of TGNNs thus effectively mixes information from the past and the future. This violates the arrow of time and questions the applicability of TGNNs in real-world prediction settings, where models do not have access to future information.
@@
Figure 1: Illustration of the issues with a batch-based evaluation of TGNNs: (a) A continuous-time temporal graph, split into batches with sizes $b=10$ (top), $b=12$ (middle), and time windows with duration $h=6$ (bottom). (b) A discrete-time temporal graph, split into batches with size $b=9$ (left), $b=10$ (middle), and time windows with duration $h=1$ (right). Splitting temporal graphs with inhomogeneous temporal activities into batches with fixed size $b$ assigns edges in time windows of varying lengths to the same batch and edges with identical timestamps to different batches. We use normalized mutual information (NMI) between the edges’ timestamps and their associated batch number (shown by colors) to quantify how much temporal information can be recovered from the sequence of batch numbers alone. In our work, we propose a time-window-oriented approach to evaluate dynamic link prediction that mitigates the information loss of current batch-based evaluation.
@@
  • We quantify the information loss due to the aggregation of edges into batches on 14 discrete- and continuous-time temporal graphs, thus showing how the dynamic link prediction task depends on the batch size.
  • To mitigate this issue, we formulate the task as link forecasting using a time window-oriented evaluation that adequately considers the available temporal information.
  • We perform an experimental evaluation of state-of-the-art TGNNs for link forecasting. Our results highlight substantial differences in model performance compared to a batch-oriented evaluation of link prediction, thus demonstrating the real-world impact of our work.
While batch-oriented processing is a technical necessity for efficient model training, our work shows that tuning the batch size essentially tunes the link prediction task, thus fitting the task to the model and undermining a fair comparison of temporal graph learning techniques. Proposing a time window-oriented evaluation of dynamic link forecasting, our work provides a simple yet effective solution to an important open issue in the evaluation of temporal graph neural networks.
@@ 2 Preliminaries and related work
A temporal (or dynamic) graph $G=\left(V,E\right)$ is a tuple where $V$ is the set of $n=\left|V\right|$ nodes and $E$ is a chronologically ordered sequence of $m=|E|$ time-stamped edges defined as $E=\left((u_{0},v_{0},t_{0}),\dots,(u_{m-1},v_{m-1},t_{m-1})\right)$ with $1\leq t_{0}\leq\dots\leq t_{m-1}\leq t_{max}$ [BetterEvaluationDynamic2022poursafaei, InductiveRepresentationLearning2021wang, BetterDynamicGraph2023yu]. Each node $v_{i}$ can have static node features $\mathbf{h_{i}}\in H_{V}$ and each edge $(u_{i},v_{j},t)$ can have edge features $\mathbf{e_{ij,t}}\in H_{E}$ that change over time. We assume that interactions occur instantaneously with discrete timestamps $t\in\mathbb{N}$. Although timestamps $t\in\mathbb{N}$ are discrete, such temporal graphs are often referred to as continuous-time [RepresentationLearningDynamic2020kazemi, FoundationsModelingDynamic2021skarding]. In contrast, discrete-time temporal graphs coarse-grain time-stamped edges into a sequence of static snapshot graphs $\{G_{t_{i}:t_{j}}\}$, where $G_{t_{i}:t_{j}}=(V,E_{t_{i}:t_{j}})$ with $E_{t_{i}:t_{j}}=\left\{(u,v)\mid\exists(u,v,t)\in E:t_{i}\leq t\leq t_{j}\right\}$ [DynamicNetworkEmbedding2022xue].
@@
Given time-stamped edges up to time $t$, the goal of dynamic link prediction is to predict whether an edge $(v,u,t+1)$ exists at future time $t+1$ [BetterDynamicGraph2023yu, BetterEvaluationDynamic2022poursafaei, RepresentationLearningDynamic2020kazemi, InductiveRepresentationLearning2021wang]. In practice, it is often computationally infeasible to train and evaluate models on all possible edges one edge at a time. Thus, the chronologically ordered sequence of edges $E$ is usually divided into temporal batches $B^{+}_{i}$, where each batch has a fixed size of $b$ edges. Edges within the same batch are typically processed in parallel [su2024pres, TemporalGraphNetworks2020rossi], thereby losing the temporal information inside each batch. In addition to the existing (positive) edges $(u,v)\in B^{+}_{i}$, non-existing (negative) edges $(u^{-},v^{-})\in B_{i}^{-}$ are sampled and used for training and evaluation. This is done since real-world graphs are typically sparse and using all possible edges between all node pairs would lead to a large class imbalance and longer runtime. With these assumptions, we can formally define the task as follows:
Definition 2.1 (Dynamic link prediction).
Let $G=(V,E)$ be a temporal graph with node features $H_{V}$ and edge features $H_{E}$. Let $b$ be the batch size and $B_{i}^{+}:=\{(u,v)\mid\exists(u,v,t)\in E\text{ with }t\in\{t_{i\cdot b},\ldots,t_{i\cdot(b+1)-1}\}\}$ the set of $b$ edges in the $i$-th batch. We further use $B_{i}^{-}$ to denote a set of negative edges drawn using negative sampling as described in Appendix A. For a given batch $i$ we use $\hat{E}_{i}=\left\{(u,v,t)\mid\exists(u,v,t)\in E:t<t_{i\cdot b}\right\}$ to denote the past edges. The goal of dynamic link prediction is to find a model $f_{\theta}\bigl{(}u,v\mid\hat{E}_{i},H_{V},H_{\hat{E}_{i}}\bigr{)}$ with parameters $\theta$ that, for each batch $i$, predicts whether $(u,v)\in B_{i}^{+}$ or $(u,v)\in B_{i}^{-}$.
@@
[…]
Temporal graph training and evaluation.
[…]
@@ 3 From link prediction to link forecasting
Learning temporal patterns in a batch-oriented fashion leads to issues in continuous-time and discrete-time graphs. Below, we show that batching leads to inconsistent tasks because the time window for prediction varies for temporal batches across different link densities in time. Temporal batches further cause information loss by either inducing a non-existing temporal order between links or ignoring the existing order. We demonstrate these issues in eight continuous-time and six discrete-time temporal graphs, whose characteristics are summarized in Table 1 and Appendix B. To mitigate these issues, we then formulate the Link Forecasting task based on fixed-length time windows.
@@
Table 1: Characteristics of continuous and discrete-time temporal graphs [BetterEvaluationDynamic2022poursafaei, BetterDynamicGraph2023yu]. For each data set, we include the type, the number of nodes $n$, the number of edges $m$, the resolution of timestamps, the total duration $T$ of the observation, the average number of edges $\overline{|E_{t}|}$ with the same timestamp $t$, and the temporal density $\nicefrac{{T}}{{m}}$.
Dataset Type $n$ $m$ Resolution $T$ $\overline{|E_{t}|}$ $\nicefrac{{T}}{{m}}$
Enron Contin. 184 125 235 1 second 3.6 years 5.5 ± 16.6 908.2 s
UCI Contin. 1899 59 835 1 second 193.7 days 1.0 ± 0.3 279.7 s
MOOC Contin. 7144 411 749 1 second 29.8 days 1.2 ± 0.5 6.2 s
Wiki. Contin. 9227 157 474 1 second 31.0 days 1.0 ± 0.2 17.0 s
LastFM Contin. 1980 1 293 103 1 second 4.3 years 1.0 ± 0.1 106.0 s
Myket Contin. 17 988 694 121 1 second 197.0 days 1.0 ± 0.0 24.5 s
Social Contin. 74 2 099 519 1 second 242.3 days 3.7 ± 2.5 10.0 s
Reddit Contin. 10 984 672 447 1 second 31.0 days 1.0 ± 0.1 4.0 s
UN V. Discrete 201 1 035 742 1 year 71.0 years 14 385.3 ± 7142.1 36.1 min
US L. Discrete 225 60 396 1 congress 11.0 congr. 5033.0 ± 92.4 1.8 $\cdot 10^{-4}$ congr.
UN Tr. Discrete 255 507 497 1 year 31.0 years 15 859.3 ± 3830.8 32.1 min
Can. P. Discrete 734 74 478 1 year 13.0 years 5319.9 ± 1740.5 91.8 min
Flights Discrete 13 169 1 927 145 1 day 121.0 days 15 796.3 ± 4278.5 5.4 s
Cont. Discrete 692 2 426 279 5 minutes 28.0 days 300.9 ± 342.4 1.0 s
Different time window durations in fixed-size batches.
One issue of batch-oriented temporal graph learning and dynamic link prediction is that activities in real-world temporal graphs are inhomogeneously distributed across time. In Figure 2 we show the temporal activity in terms of the number of time-stamped edges within a given time interval both for continuous-time and discrete-time temporal graphs. For continuous-time data, we used binning in six-hour intervals. The results show that most real-world temporal graphs have highly inhomogeneous activities across time. For batch-oriented evaluation, this introduces the issue that each fixed-size batch $B_{i}^{+}$ determines a time window with duration $t_{j}-t_{i}$ that is shorter (longer) during periods with higher (lower) activity.
In Figure 3 we evaluate the dependency between batch size and window for empirical temporal graphs. We observe that, both in continuous- and discrete-time temporal graphs, a single batch size can create time windows with varying durations even within the same dataset. For continuous-time temporal graphs, we typically have much bigger batches than edges per timestamp such that the time window of a batch is long (cf. Table 1). The number of edges per snapshot in discrete-time temporal graphs is generally larger than the batch size $b$ in any period regardless of the density (Table 1). This means that edges in a batch often belong to the same snapshot leading to small window durations.
(a) Continuous-time temporal graphs resoluted into seconds and binned into 6-hour time periods.
(b) Discrete-time temporal graphs with different time resolution (see Table 1).
Figure 2: Real-world datasets exhibit diverse edge occurrence patterns that are visualised using the edge density across time, i.e., histograms counting the number of edges per timestamp. Dashed lines divide the datasets into train, validation, and test sets as used in Section 4.
As an example, consider the Myket dataset [EffectChoosingLoss2023loghmania] which contains users $v$ and Android applications $u$, connected at time $t$ when user $v$ installs application $u$. The timestamps are provided in seconds and edges occur roughly every half a minute on average (cf. Table 1), making the expected time range for a batch with size $b=2$ approximately 0.5 minutes. With $b=2$, the task is to predict which users install what applications during this time window. Choosing $b=120$ or $b=2880$ turns the task into a prediction problem for approximately the next hour or day, respectively. As we can see, batching not only leads to incomparable prediction tasks between models and datasets due to the varying window duration but also acts as a kind of coarse-graining discarding all temporal information inside each batch. We discuss this information loss and resulting issues in the next section.
Figure 3: Scatterplots visualising time window durations of all batches of size $b$ for different $b$s.
Information loss in batch-based temporal graph learning.

### v2

# From Link Prediction to Forecasting: Addressing Challenges in Batch-based Temporal Graph Learning

@@ Abstract
Dynamic link prediction is an important problem often considered in recent works proposing various approaches for learning temporal edge patterns. To assess their efficacy, models are evaluated on benchmark datasets involving continuous-time and discrete-time temporal graphs. However, as we show in this work, the suitability of common batch-oriented evaluation depends on the datasets’ characteristics, which can cause multiple issues: For continuous-time temporal graphs, fixed-size batches create time windows with different durations, resulting in an inconsistent dynamic link prediction task. For discrete-time temporal graphs, the sequence of batches can additionally introduce temporal dependencies that are not present in the data. In this work, we empirically show that this common evaluation approach leads to skewed model performance and hinders the fair comparison of methods. We mitigate this problem by reformulating dynamic link prediction as a link forecasting task that better accounts for temporal information present in the data.
Graph Neural Network, GNN, Temporal Graph, Dynamic Link Prediction, Dynamic Graph, Temporal Graph Learning, Dynamic Graph Learning, Temporal Graph Neural Network, TGNN, DyGNN, Dynamic Graph Neural Network
@@ 1 Introduction
Figure 1: Illustration of the issues in batch-based evaluation: (1) Information loss: continuous-time temporal edges are grouped into batches, causing some TGNNs to discard their temporal ordering. (2) Information leakage: discrete-time temporal edges that belong to the same snapshot are assigned to different batches, imposing an artificial ordering between edges that belong to the same snapshot. (3) Varying time window durations: using a fixed batch size creates time windows of different lengths when temporal edges are inhomogeneously distributed. (4) Tunable batch size: when different models choose different batch sizes, they group the temporal edges into different time windows, thus making their results incomparable because the specifics of their prediction tasks differ.
Many scientific fields study data that can be modelled as graphs, where nodes represent entities that are connected by edges. Examples include social (Lazer et al., 2009), financial (Bardoscia et al., 2021), as well as molecular networks (David et al., 2020). Apart from the mere topology of interactions, i.e., who is connected to whom, such data increasingly include information on when these interactions occurred. Depending on the temporal resolution, the resulting temporal graphs are often categorised as continuous-time or discrete-time (Kazemi et al., 2020): State-of-the-art data collection technology provides high-resolution continuous-time temporal graphs that capture each interaction’s exact (and possibly unique) occurrence time, leading to temporal graphs that, e.g., capture time-stamped online interactions (Kumar et al., 2019) or contacts between individuals via proximity (Vanhems et al., 2013). In contrast, discrete-time temporal graphs give rise to a temporally ordered sequence of static snapshots, where each snapshot contains interactions recorded within a (typically coarse-grained) time interval. Examples include political collaboration networks (Fowler, 2006) frequently available in yearly snapshots.
Building on the growing importance of temporal data and the success of graph neural networks (GNNs) for static graphs (Corso et al., 2024), deep graph learning has recently been extended to temporal (or dynamic) graphs (Feng et al., 2024). To this end, several temporal graph neural network (TGNN) architectures have been proposed that are able to simultaneously learn temporal and topological patterns. These architectures are often evaluated in dynamic link prediction, where the task is to predict the existence of edges during a future time window of length $\Delta t$, e.g., to provide recommendations to users (Kumar et al., 2019).
TGNNs utilise temporal batches to speed up training for dynamic link prediction. To construct these temporal batches, the sequence of temporally ordered edges is divided into a sequence of equally large chunks, i.e. containing the same number of edges, leading to several problems (cf. Figure 1): (1) Some TGNNs treat edges within a batch as if they occurred simultaneously, essentially discarding temporal information within a batch for continuous-time data. (2) In discrete-time temporal graphs, large snapshots containing edges with identical timestamps are often split into a sequence of batches. This introduces a temporal order not present in the data, thus leaking information that should not be available to the model because edges with the same timestamp end up in different batches. (3) Fixed-size batches in continuous-time temporal graphs are often associated with time windows of varying lengths ${\Delta t}$, effectively resulting in tasks of varying difficulty for each batch. (4) Related to the previous issue, treating the batch size as a tunable parameter affects the resulting window lengths and changes the task difficulty, e.g., from predicting at the minute to the hour level, thus making model performance between different models incomparable when they choose different batch sizes.
@@
  • We quantify the information loss and leakage due to the aggregation of edges into batches on 14 discrete-time and continuous-time temporal graphs showing how dynamic link prediction performance can be tuned by changing the batch size.
  • Replacing the current link prediction task, we formulate a link forecasting task based on time windows instead of batches to adequately consider the available temporal information.
  • We experimentally evaluate state-of-the-art TGNNs for link forecasting. Our results highlight substantial differences in model performance compared to the batch-based link prediction task. Our results suggest that memory-based TGNNs are ill-suited for discrete-time data, which has so far been overlooked due to the overestimation of model performance caused by information leakage in batch-based evaluation.
  • We provide implementations of the link forecasting task for commonly used graph learning frameworks such as DyGLib (Yu et al., 2023) and PyTorch Geometric (Fey & Lenssen, 2019).
Although batch-based processing is a technical necessity for efficient model training, we show that splitting edges into batches regardless of their occurrence time leads to many problems. With dynamic link forecasting, we propose a simple yet effective solution. While having similar runtime as dynamic link prediction, our time-window-based evaluation facilitates a fairer and more realistic evaluation that better reflects real-world scenarios.
@@ 2 Preliminaries and Related Work
A temporal, or dynamic graph $G=\left(V,E\right)$ is a tuple where $V$ is the set of $n=\left|V\right|$ nodes and $E$ is a chronologically ordered sequence of $m=|E|$ time-stamped edges defined as $E=\left\{(u_{1},v_{1},t_{1}),\dots,(u_{m},v_{m},t_{m})\right\}$ with $t_{1}\leq\dots\leq t_{m}$ (Wang et al., 2021b; Yu et al., 2023). Each node $v_{i}$ can have static node features $\mathbf{h_{i}}\in H_{V}$ and each edge $(u_{i},v_{j},t)$ can have edge features $\mathbf{e_{ij,t}}\in H_{E}$ that change over time. We assume that interactions occur instantaneously with timestamps $t\in\mathbb{N}$. Although timestamps $t\in\mathbb{N}$ are discrete, such temporal graphs are often categorized as continuous-time (Kazemi et al., 2020; Skarding et al., 2021). In contrast, discrete-time temporal graphs coarse-grain time-stamped edges into a sequence of static snapshots $\left(G_{t_{i}:t_{j}}\right)$, where $G_{t_{i}:t_{j}}=(V,E_{t_{i}:t_{j}})$ with $E_{t_{i}:t_{j}}=\left\{(u,v)\mid\exists(u,v,t)\in E:t_{i}\leq t\leq t_{j}\right\}$ (Xue et al., 2022).
@@
Given time-stamped edges up to time $t$, the goal of dynamic link prediction is to predict whether an edge $(v,u,t+1)$ exists at future time $t+1$ (Kazemi et al., 2020; Wang et al., 2021b; Poursafaei et al., 2022; Yu et al., 2023). In practice, it is often computationally infeasible to train and evaluate models on all possible edges one edge at a time. Thus, the chronologically ordered sequence of edges $E$ is usually divided into temporal batches $B^{+}_{i}$, where each batch has a fixed size of $b$ edges. Edges within the same batch are typically processed in parallel (Rossi et al., 2020; Su et al., 2024). In addition to the existing (positive) edges $(u,v,t)\in B^{+}_{i}$, non-existing (negative) edges $(u^{-},v^{-},t^{-})\in B_{i}^{-}$ are sampled and used for training and evaluation (see Appendix A for a brief overview of negative sampling approaches). This is done since real-world graphs are typically sparse and using all possible edges between all node pairs would lead to a large class imbalance and longer runtime. Formally, we define the dynamic link prediction task as follows:
Definition 2.1.
Let $G=(V,E)$ be a temporal graph with node features $H_{V}$ and edge features $H_{E}$. Let $b$ be the batch size and $B_{i}^{+}:=\{(u_{j},v_{j},t_{j})\mid\exists(u_{j},v_{j},t_{j})\in E:b\cdot i\leq j<b\cdot(i+1)\}\}$ the set of $b$ edges in the $i$-th batch. We further use $B_{i}^{-}$ to denote a set of negative edges drawn using negative sampling as described in Appendix A. For a given batch $i$ we use $\hat{E}_{i}=\left\{(u,v,t)\mid\exists(u,v,t)\in E:t<t_{b\cdot i}\right\}$ to denote the past edges. The goal of dynamic link prediction is to find a model $f_{\theta}\bigl{(}u,v,t\mid\hat{E}_{i},H_{V},H_{\hat{E}_{i}}\bigr{)}$ with parameters $\theta$ that, for all edges $(u,v,t)\in B_{i}^{+}\cup B_{i}^{-}$ in each batch $i$, predicts whether $(u,v,t)\in B_{i}^{+}$ or $(u,v,t)\in B_{i}^{-}$.
@@
Current state-of-the-art dynamic link prediction methods, such as JODIE (Kumar et al., 2019), DyRep (Trivedi et al., 2019), TGN (Rossi et al., 2020) keep an up-to-date memory of temporal information in the graph by utilizing recurrent neural networks. TGAT extends graph attention to the temporal domain and replaces positional encodings in GAT (Velickovic et al., 2018) with a vector representation of time (Xu et al., 2020). TCL (Wang et al., 2021a) uses a transformer-based architecture to capture the nodes’ time-evolving properties. CAWN learns temporal motifs based on causal anonymous walks (Wang et al., 2021b). GraphMixer takes an attention-free and transformer-free approach, using an MLP-based link encoder, a mean-pooling-based node encoder, and an MLP-based link classifier for predictions (Cong et al., 2023). DyGFormer combines nodes’ historical co-occurrences as interaction targets of the same source node with a temporal patching approach to capture long-term histories (Yu et al., 2023). Further approaches for temporal graph learning exist, including DyGEM (Taheri et al., 2019), DySAT (Sankar et al., 2020), EvolveGCN (Pareja et al., 2020), and DBGNN (Qarkaxhija et al., 2022). For a recent survey of deep-learning-based dynamic link prediction, we refer to Feng et al. (2024).
Open Challenges in Dynamic Link Prediction
Training.
Recent works (Zhou et al., 2022, 2023; Feldman & Baskin, 2024; Su et al., 2024) identified issues in the training of memory-based TGNNs with large batch sizes: Processing edges belonging to the same batch in parallel ignores their temporal dependencies, resulting in varying performance depending on the chosen batch size. This issue, called temporal discontinuity, is caused by Problem (1) from above, i.e. treating edges within the same batch as if they occurred simultaneously (cf. Appendix D). To solve this issue, Zhou et al. (2023) introduce a distributed training framework using smaller batch sizes. The proposed solution PRES (Su et al., 2024) accounts for intra-batch temporal dependencies using a prediction-correction scheme. Feldman & Baskin (2024) propose a model with a decoupling strategy to mitigate this problem. However, these works focus on training instead of evaluation which is our work’s focus. Parallel to our work, Huang et al. (2024) present the unified temporal graph (UTG) representation, which facilitates the training of continuous-time graphs with discrete-time TGNNs. They also train continuous-time TGNNs on discrete-time graphs and, thereby, briefly mention information leakage as described in Problem (2). However, this work does not consider all Problems (1) – (4), nor does it include an empirical evaluation of how Problem (2) affects model performance, which is addressed in our work.
Evaluation.
Recent progress in terms of TGNN evaluation includes the temporal graph benchmark (TGB) (Huang et al., 2023) similar to the static open graph benchmark (OGB) (Hu et al., 2020). Poursafaei et al. (2022) identify problems with random negative sampling for dynamic link prediction and propose novel time-dependent negative sampling techniques to improve the evaluation of TGNNs. Gastinger et al. (2023) identify issues in the evaluation of temporal knowledge graph forecasting. Although none of the models used for this task overlap with regular TGNNs for dynamic link prediction, some of the problems can be related, e.g., differences in forecasting horizons leading to incomparable results.
Research Gap
The aforementioned issues in training and evaluating TGNNs for dynamic link prediction arise from applying standard procedures from static link prediction to temporal graph data. While recent works have identified how some of these issues affect the performance of some models in some data sets, we lack a systematic understanding of how the Problems (1) - (4) affect state-of-the-art temporal graph learning architectures across different types of continuous- and discrete-time temporal graph data. Moreover, we lack a better formulation of a dynamic link forecasting task, which mitigates these problems and can guide the development of new TGNNs.
@@ 3 From Link Prediction to Link Forecasting
Learning temporal patterns in a batch-oriented fashion leads to issues in continuous-time and discrete-time graphs. We demonstrate these issues in eight continuous-time and six discrete-time temporal graphs, whose characteristics are summarised in Table 1 and Appendix C. To mitigate these issues, we formulate the link forecasting task based on time windows with fixed duration.
@@
Table 1: Characteristics of continuous (top) and discrete-time (bottom) temporal graphs. For each dataset, we list the number of nodes $n$ and edges $m$, the resolution of timestamps, the dataset’s total duration $T$, the average number of edges $\overline{|E_{t}|}$ with the same timestamp $t$, and the temporal density $\nicefrac{{T}}{{m}}$. The US Legis. dataset is measured in congresses (c). The other datasets are measured in conventional units of time,i.e. seconds (s), minutes (min), days (d) or years (a). See Appendix C for more information on each dataset.
Dataset $n$ $m$ Res. $T$ $\overline{|E_{t}|}$ $\nicefrac{{T}}{{m}}$
Enron 184 125 235 1 s 3.6 a 5.5 ± 16.6 908.2 s
UCI 1899 59 835 1 s 193.7 d 1.0 ± 0.3 279.7 s
MOOC. 7144 411 749 1 s 29.8 d 1.2 ± 0.5 6.2 s
Wiki. 9227 157 474 1 s 31.0 d 1.0 ± 0.2 17.0 s
LastFM 1980 1 293 103 1 s 4.3 a 1.0 ± 0.1 106.0 s
Myket 17 988 694 121 1 s 197.0 d 1.0 ± 0.0 24.5 s
Social 74 2 099 519 1 s 242.3 d 3.7 ± 2.5 10.0 s
Reddit 10 984 672 447 1 s 31.0 d 1.0 ± 0.1 4.0 s
UN V. 201 1 035 742 1 a 71.0 a 14 385.3 ± 7142.1 36.1 min
US L. 225 60 396 1 c 11.0 c 5033.0 ± 92.4 1.8 $\cdot 10^{-4}$ c
UN Tr. 255 507 497 1 a 31.0 a 15 859.3 ± 3830.8 32.1 min
Can. P. 734 74 478 1 a 13.0 a 5319.9 ± 1740.5 91.8 min
Flights 13 169 1 927 145 1 d 121.0 d 15 796.3 ± 4278.5 5.4 s
Cont. 692 2 426 279 5 min 28.0 d 300.9 ± 342.4 1.0 s
3.1 Problems in Batch-based Dynamic Link Prediction
Below, we show that batching causes information loss or leakage by ignoring the existing temporal order of links or inducing an artificial order between them, respectively. Temporal batches lead to inconsistent tasks because they produce prediction time windows of varying lengths when link densities vary over time. The batch-based evaluation can further lead to incomparable results between models when the batch size is treated as a hyperparameter and different models choose different batch sizes.
(1) Information Loss.